### v1

# Paired Completion: Flexible Quantification of Issue-framing at Scale with LLMs

Abstract
Detecting and quantifying issue framing in textual discourse - the perspective one takes to a given topic (e.g. climate science vs. denialism, misogyny vs. gender equality) - is highly valuable to a range of end-users from social and political scientists to program evaluators and policy analysts. However, conceptual framing is notoriously challenging for automated natural language processing (NLP) methods since the words and phrases used by either ‘side’ of an issue are often held in common, with only subtle stylistic flourishes separating their use. Here we develop and rigorously evaluate new detection methods for issue framing and narrative analysis within large text datasets. By introducing a novel application of next-token log probabilities derived from generative large language models (LLMs) we show that issue framing can be reliably and efficiently detected in large corpora with only a few examples of either perspective on a given issue, a method we call ‘paired completion’. Through 192 independent experiments over three novel, synthetic datasets, we evaluate paired completion against prompt-based LLM methods and labelled methods using traditional NLP and recent LLM contextual embeddings. We additionally conduct a cost-based analysis to mark out the feasible set of performant methods at production-level scales, and a model bias analysis. Together, our work demonstrates a feasible path to scalable, accurate and low-bias issue-framing in large corpora.
1 Introduction
It is widely held that public narratives have the power – for better and worse – to shape society (Shiller 2019; Patterson and Monroe 1998; Graber 2002; Barabas and Jerit 2009). For quantitative social scientists, a typical analytical strategy to quantify the occurrence, characteristics and dynamics of these important narratives is to use a ‘framing’ lens. According to the much-cited definition found in Entman (1993), framing is the process by which individuals “select some aspects of a perceived reality and make them more salient in a communicating text” with the purpose of promoting a particular interpretation or evaluation of reality. In essence, to frame, is to impose a world-view or ‘way-of-thinking’ in communication, with the hope that others will be persuaded to be convinced of the same. In the standard approach to framing analysis (Chong and Druckman 2007), one begins first, by identifying an issue (e.g. ‘climate change’); second, by defining the dimensions of that issue (e.g. ‘causes’, ‘economic impact’, etc.); third, by developing framings of those dimensions (e.g. climate change/ causes/ framing: ‘anthropogenic emissions are responsible for climate change’); and then finally, by the manual labelling of texts (sentences, paragraphs) as to their framing alignment. Computational approaches to framing quantification have addressed various aspects of this decomposition, often (unhelpfully) under the generic heading of ‘framing’ (we return to this point below) (Ali and Hassan 2022).
Almost all prior approaches to computational framing analysis consider the task as a supervised machine-learning problem, typically as a multi-class classification task (Field et al. 2018), and most often focusing on automatic labelling of dimensions (e.g. ‘2nd Amendment’, ‘Politics’, ‘Public Opinion’) of a single issue (e.g. gun violence) (Liu et al. 2019a; Zhang et al. 2023), rather the more elaborate ‘world-view’ like conceptualisation that Entman (1993) and Chong and Druckman (2007) hold. Where studies consider conceptual framing identification, large amounts of labelled data are required, and reported accuracy is very low (below 0.6) (Morstatter et al. 2018; Mendelsohn, Budak, and Jurgens 2021), demonstrating the severe challenges inherent in automating an already difficult human-level task.
Figure 1: Paired Completion – Target texts ($X$) are taken, one at a time, as completions to one ($k=1$) or two ($k=2$) priming conditioner sentences from two opposing issue framing sets, $A$, $B$, in turn. Summed log-probabilities of each completion text ($x$) are obtained from the LLM as if the LLM had used the text to follow the conditioners. The two resultant summed log-probs inform the $Delta$ metric.
In this study, we introduce paired completion – a low-resource, ‘few-label’, computationally efficient method that can accurately identify whether a target text aligns with one or other conceptual framing on a given issue (see Fig. 1). Importantly, and distinguishing it from previous methods, our approach: requires only a few (e.g. 5-10) example texts of a given framing (in fact, these can be generatively created); is low-bias compared to generative (prompt-based) LLM approaches; is significantly cheaper than generative approaches; and is highly flexible, switching issues or dimensions or framings is trivial.
Paired completion takes advantage of the log-probability (logprob) outputs of an LLM [^1] Note: logprobs are available as outputs through the OpenAI API for “babbage-002” and “davinci-002” (OpenAI 2024a), and can be gathered by running the “vLLM OpenAI-compatible API” (Kwon et al. 2023) on a local machine, for a wide variety of open-source models. to find conditional probabilities of a text given a series of conditioners from different conditioning sets. We use the relative differences in probabilities to establish a baseline metric that (at least theoretically) is resilient to the model’s prior probabilities of both the conditioning text and the text being aligned to the conditioning sets. We demonstrate empirically that this method is successful, and that one achieves superior performance from using this method with raw base models compared to “asking” instruct-fine-tuned AI-models the “question” at hand.
We conduct rigorous evaluation of our proposed method across 192 independent experiments by comparing it to four framing classification approaches over four diverse, synthetic textual datasets, including two baseline approaches (traditional tf-idf vectors (Sparck Jones 1972; Salton 1983), and fasttext sentence embeddings (Bojanowski et al. 2017)) and three LLM-based methods (contextual embeddings (Peters et al. 2018; Devlin et al. 2018; OpenAI 2024b) LLM chat token probabilities (Radford et al. 2019), and our novel paired completion method). We demonstrate that the LLM-based approaches are, in general, far superior to the alternatives. The LLM-embedding approach is powerful with enough training data, but with small amounts of data (e.g. five sentences for each conditioning set) the LLM methods easily outperform LLM-embeddings. We also demonstrate that paired completion with LLMs is generally superior to the LLM prompting approach. We discuss why this might be the case in Section 2.1, from a theoretical perspective. We also conduct cost- and bias- comparison analysese at current gated API pricing to assess any trade-offs in performance.
1.1 Contributions
We introduce paired completion as a solution to the problem of textual alignment. We construct a series of high-quality synthetic datasets using a novel method which captures nuances of discourse on complex topics, and use these datasets to evaluate the performance of several approaches to textual alignment. We demonstrate that paired completion is a novel, efficient, and more effective method for performing textual alignment (compared to a chat-based LLM baseline).
1.2 Related Literature
‘Framing’ analysis
Unfortunately, ‘framing’ analysis does not have a clear definition in the computational literature, as evidenced by the variety of tasks that arise in a recent survey of 37 ‘framing’ studies (Ali and Hassan 2022). A starting point for many framing approaches is to leverage existing corpora of labelled datasets. Here, the media frames corpus (MFC) (Card et al. 2015) and the gun violence frame corpus (GVFC) (Liu et al. 2019a) have been the basis of many methodological contributions. However, these datasets conceptualise ‘framing’ as dimensions (ala Chong and Druckman (2007)) of a topic or issue, not conceptual frames as we have distinguished earlier. The MFC is composed of 15 generic ‘frames’ (Boydstun et al. 2013) such as ‘economic’, ‘public opinion’ and ‘cultural identity’ applied to three issues (‘immigration’, ‘smoking’, and ‘same-sex marriage’). Thousands of annotations were recorded as to whether one of these dimensions were associated with the text on a given issue. Likewise, the GVFC follows a similar approach, albeit tied more tightly to gun violence, ‘frames’ are equivalent to dimensions, and include ‘2nd Amendment’, ‘Politics’, and ‘Public Opinion’. Typically, computational methods approach framing in this way as a multi-class classification problem, using supervised machine learning methodologies such as featuring engineering, classifier selection and k-fold evaluation (Field et al. 2018; Liu et al. 2019a; Akyürek et al. 2020; Zhang et al. 2023). Common to all of these approaches is the need for large amounts of labelled ground-truth data, and consequently, the outcome that methods are not generalisable beyond the topics under study.
Where ‘framing’ is implemented in a closer way to the conceptual framing we address in this work, challenges remain around the need for large labelled datasets and the accuracy of the methods. Morstatter et al. (2018) consider support for, or against, 10 framings related to Balistic Missile Defence (BMD) in Europe over 823 online news articles (31k sentences). By writing out support- and opposition- (polarity) perspectives for each of the 10 framings, they are able to generate 20 framing-polarity classes, and apply traditional NLP methods to multi-class prediction. Alternatively, Mendelsohn et al. (2021) label 3.6k social media posts (tweets) related to immigration for the 15 generic ‘framings’ of Boydstun et al. (Boydstun et al. 2013) together with 11 conceptual framings (e.g. ‘hero’, ‘victim’, ‘threat’ positions on immigrants). Using a base and fine-tuned encoder-only transformer approach (Devlin et al. 2018; Liu et al. 2019b), they again conduct a multi-label classification study. Whilst these examples are closely aligned to the same conceptual framing identification problem we address in the current study, each requires thousands of hand-labelled data to develop features to train traditional supervised machine-learning algorithms. Furthermore, and underlining the challenge of this task for traditional (even transformer based encoder methods), accuracy scores across 20 polarity classes in Morstatter et al. and f1 scores across 11 conceptual framings in Mendelsohn, Budak, and Jurgens are up to just 0.434 and 0.552 respectively.
Alternatively, and closer to the spirit of the present work, Guo, Ma, and Vosoughi (2022) take on a related but distinct task of quantifying the similarity between news sources by fine-tuning LLMs to each source and then conducting differential experiments on the likelihoods of word-completions when masking specific words within sentences on specific topics. The approach is highly interesting and would enable the characterisation of how a source (e.g. a news outlet, a speaker, a social media account) conceptually framed a specific policy area relative to another source. However, fine-tuning a LLM requires a large amount of text from each source to be feasible, more than would be available for more than just the most prominent political- or business- leaders. And, whilst the method could be applied to sub-sets of the corpora by time-interval (reminiscent of computational cultural analysis (Kozlowski, Taddy, and Evans 2019)) this further reduces the texts available.
Together, existing computational approaches to ‘framing’, whether focusing on the different, higher-level task of dimension identification, or studying conceptual framing as we do in the present work, require large amounts of labelled data to train multi-class classifiers. In contrast, paired completion, studies in the present work, proceeds with minimal examples of the conceptual framings required (e.g. 5-10) and requires no ‘training’ as such.
Stance detection
Finally, a related task is that of stance detection, which we address briefly. Stance detection is typically formalised as the, “automatic classification of the stance of the producer of a piece of text, towards a target, into one of three classes: Favor, Against, Neither.” (Küçük and Fazli 2020) In effect, stance detection is a sub-problem of sentiment analysis, and again, typical approaches leverage traditional NLP techniques with labelled data as inputs (Küçük and Fazli 2020). Whilst paired completion shares the notion of ‘target texts’, these texts work in concert to mark out a complex, nuanced conceptual framing on any issue, and the two priming sets (A,B) need not be strictly in opposition, but represent two perspectives, opening up more complex analytical insights than simply ‘favor’ or ‘against’.
Perplexity
One common measure of the capability of an LLM is perplexity (Jelinek et al. 1977), which is a statistical measure of the model’s “surprise” at a given completion under the logic that a model which is less surprised by correct answers is better (similar to the maximum likelihood principle). The paired completion approach developed in this work is a measure similar to perplexity, but instead of seeking the estimated likelihood of a particular completion we instead calculate and compare the likelihoods of multiple completions of a given text.
2 Textual Alignment & Paired Completion
To hone in on Entman’s classic definition of framing, we reconceptualise the problem as one of “textual alignment”. Namely, two texts on some topic or issue arise from the same conceptual framing, if they share a high level of textual alignment – a measure of the likelihood (in some sense) that the two texts might be spoken by the same entity (with a constant conceptual framing). This implies the statements come from the same theoretical outlook, model of the world, and/or causal structure. It is important that the expressive entity is generally defined. For we will be, at times, leveraging generative AI LLMs to play the role of $\mathfrak{E}$, alongside human expression, to quantify the degree of alignment.
Definition 1 (Textual alignment).
Given two conditioning texts $a$ and $b$, and an expressive entity, $\mathfrak{E}$ (e.g. a person, a generative AI LLM), text $x$ is said to be more textually aligned with $a$ versus $b$ if it is more likely that $x$ would be expressed by some $\mathfrak{E}^{\prime}$ who previously expressed $a$, than the alternate case where $\mathfrak{E}^{\prime}$ had previously expressed $b$.
Importantly, Def 1 is not the same as similarity. Consider the texts, ‘Getting a dog will improve your life’ and, ‘Getting a dog will ruin your life’. Whilst these are very similar (in fact, an LLM-powered contextual similarity score would be close to 1 for these texts), they are not textually aligned. If someone holds the view that dogs improve your life (framing A), it is highly unlikely that they would say that dogs ruin your life (framing B). Yet these texts are highly similar on sentiment (both are neutrally posed) and share an almost identical vocabulary. However, consider the third text, ‘Pets help to keep you fit and healthy’. It is clear that this text is strongly textually aligned with framing A, but strongly dis-aligned with framing B. Yet, this text is perfectly dissimilar in vocabulary, and is of middling similarity in an LLM-powered contextual embedding space. These examples demonstrate that issue-framing, formalised as textual alignment, is both ‘simple’ for a human to perceive, yet difficult for existing computational methods (based on similarity, sentiment, vocab, embeddings) to detect.
As such, we desire a new set of tools to quantify textual alignment. We consider these tools in the context of the “Issue-Framing” task, where a user wishes to detect and quantify texts from a large corpus which share the same framing, via textual alignment. Suppose the user has a small set of texts which together lay out a given framing position A, as compared to an opposing framing position B with a similar number of texts. We then formalise this task as follows:
Definition 2 (The Issue-Framing Task).
Given a corpus of texts $X$ (target texts) and a set of priming (or framing) texts $S=\{A,B\}$ comprising texts which represent framing A and B, for each $x\in X$, quantify the textual alignment towards $A$ and $B$.
Naturally, the user could accomplish this task by reading every text in $X$ and marking (labelling) whether the text is textually aligned with the conditioning or framing texts from $A$ or $B$. However, the aim of our work is to develop methods that might reliably accomplish this task at scale in an automated manner.
We conjecture that LLMs are well suited to performing the issue-framing task since, with the advent of attentional transformer technology, they have been shown to be remarkably successful at modelling human language. That is, forming a deep abstract representation of the meaning of human communication. Precisely the capability we require to assess textual alignment.
2.1 Paired Completion
We propose the “paired completion” method as a solution for the textual-alignment definition given above. Figure 1 gives an overview of its components. Given some set of target texts on a given topic we wish to analyse, and a small set of texts which provide frames for perspective A and B on a given topic (e.g. ‘get a dog’ vs. ‘don’t get a dog’), we construct a pair of prompt sequences, $s_{1}+x$ and $s_{2}+x$ to pass to a generative LLM. Each prompt sequence is composed of a random selection from one of the priming sets (e.g. $s_{1}$ ‘get a dog’), followed by the target text ($x$).
For example, a prompt sequence could be ‘[priming text from A, $s_{1}$] Owning a dog will improve your life. [target text, $x$] Dog owners have lower blood pressure and less stress in general.’ A similar sequence would be created for the same text $x$ with priming text(s) from set B. Each prompt sequence is then passed, one at a time, to a generative LLM, and instead of seeking a completion (i.e. generating new tokens) from the LLM, we instead exploit many LLM’s ability to provide log-probabilities (the log of the likelihood that the model would have chosen that token/word next) for each token passed to the language model as if it had generated this exact sequence of text. By so doing, we generate two conditional log-probabilities, $lp(x|s_{1})$ and $lp(x|s_{2})$ (see details in sec 2.1), the conditional log-probs of $x$ being the completion to the priming sequence $s_{1}$ and $s_{2}$ respectively.
In this way, we are leveraging the twin features of LLMs: first, that LLM attentional mechanisms are highly adept at representing the latent semantic state of a given text; and second, that LLMs have been trained to provide coherent sequences of text (i.e. to avoid non sequiturs). Together, the priming sequence will set the LLM on a particular statistical trajectory to keep the framing state consistent, which implies that if $x$ is within this trajectory (i.e. $x$ is textually aligned with the priming state), the summed log-probabilities the LLM assigns to the words in $x$ will be high. Whereas, if $x$ appears to contradict or speak for a different framing than the priming sequence, the log-probabilities for the words in $x$ will be very low. It is this difference that we exploit by testing both priming sequences from A and B to then calculate the Diff metric.
To summarise, paired-completion leverages the ‘deep’ language modelling properties of LLM – their deep contextual representation of human meaning – to quantify the likelihood that a target text will follow from a given conditioning prior. As such, we conjecture that powerful LLMs that are neither fine-tuned (to a particular task or corpus) nor moderated by post-training methods to suppress some behaviours and up-promote others (e.g. reinforcement learning with human feedback, RHLF (Ziegler et al. 2020)) will be most well suited to the paired completion method. Either of these adaptations of LLMs could reasonably trade off fundamental (general) language modelling capabilities of LLMs in service of performance or safety on a particular task or in a particular context.
See the Supplementary Information for details of the implementation of this method in evaluation.
The Diff Metric
Suppose we have a set of $n$ priming sequences, $S=\{s_{1},s_{2},...,s_{n}\}$, and a set of $m$ target sequences $X=\{x_{1},...,x_{m}\}$. We wish to find the relative alignment, in some sense, of the elements within $X$ towards the different priming sequences in $S$.
We define the diff metric as follows:
$\Delta(s_{1},s_{2},x)=\text{lp}(x|s_{1})-\text{lp}(x|s_{2})$
Note that $\Delta(s_{1},s_{2},x)=-\Delta(s_{2},s_{1},x)$.
The diff metric $\Delta$ describes the difference between the conditional probability of sentence $x$ after priming sequence $s_{1}$ and the conditional probability of sentence $x$ after priming sequence $s_{2}$. In practice, we calculate the prior probabilities of all priming sequences $s\in S$ as $p_{s}$, and all texts in $x\in X$ as $p_{x}$, and the probability of a concatenated string $s+x$ as $p_{s}x$. Note that concatenation is not necessarily simple string concatenation, but rather ensures grammatical correctness - there is no perfect way to do this, but we found that just ensuring grammatical correctness seems to work sufficiently well in practice.
We then compute $\text{lp}(x|s)=p_{sx}-p_{s}$ to find the conditional probability of $x$. We can compare this to the prior probability $p_{x}$ to determine whether the presence of $s$ has made $x$ more or less likely, and we can compute $\text{lp}(x|s_{1})-\text{lp}(x|s_{2})$ (i.e. the $\Delta$ metric). Since a larger logprob indicates a higher probability, $\Delta$ will be positive if $x$ is more likely after $s_{1}$ than after $s_{2}$, and negative if $x$ is less likely after $s_{1}$ than after $s_{2}$. Because LLMs (and language models in general) might assign different prior probabilities to both the conditioning sentences $s$ and the alignment text $x$, any such method must be robust to priors. This is why we use the difference in conditional probabilities of the same text with different prompts, which is robust to the prior probabilities of both $s$ and $x$.
One interpretation of this approach, with reference to Def. 1, is that the LLM performs the role of the expressive entity $\mathfrak{E}$, and so provides a quantification of the likelihood that the text $x$ follows text $s_{1}$, versus following text $s_{2}$, i.e. we obtain a measure of textual alignment.
Since the core idea of paired completion is to use the priming/conditioning sequence to statistically deflect the LLM towards the given framing (and so, measure the model’s degree of ‘surprise’ with the completion text) we conjecture that a longer conditioning sequence may lead to improved accuracy in classifying and retrieving texts aligned with a given framing. To explore this possibility we test two treatments, with either one ($k=1$) or two ($k=2$) priming/conditioning text(s) being used. Implementation details are provided in Supplementary Information.
3 Synthetic Dataset Formation
As described earlier, existing ‘framing’ datasets, such as the MFC (Card et al. 2015) and GVFC (Liu et al. 2019a) are not well suited for application to the task we study here since they label dimensions of a topic as ‘frames’. We are not aware of another comparable dataset that makes labels of conceptual framing available across a number of issues. For this reason, conducting evaluation with synthetic data was considered for the present study, although acknowledging inherent limitations (see last section). That said, there are some positive attributes of using synthetic data that we briefly outline.
First, our initial experience with practitioners in fields that are attempting to change public narratives, demonstrated that non-synthetic (human authored) examples of framings can carry correlated linguistic features that may pollute analysis. For example, found narratives that carry a misogynistic perspective can be relative short and abrupt, whilst opposing narratives which speak for gender equality often are expressed with longer, more complex reasoning. Early testing showed that LLMs could pick up on linguistic features such as length and complexity, confusing the signal. Whereas, our paired synthetic pipeline (see full prompts etc. in Supplementary Information) is designed to provide a very balanced (tone, length, complexity etc.) dataset, with only the conceptual framing as the distinguishing feature of the texts.
Second, we were concerned that found text could be part of the training data of the LLMs we employed (either with prompting or paired completion). By using synthetic data, although we are in-effect ‘re-generating’ realistic data, and we cannot exclude the possibility that sequences of real text were created, by using a higher temperature in generation (0.5) we are able to somewhat mitigate this. The idea being that again, our LLM methods focus on framing anlaysis, rather than ‘familiarity’. We return to this point in limitations.
The synthetic dataset generation pipeline takes a topic (e.g. “dog ownership”, “climate change”, etc.) and produces a corpus of sentences that reflect different perspectives on the topic. The generation process is a two-step hierarchical process where we generate seed perspective and then generate sentences that align with each perspective. We also generate distillations (into a smaller number of sentences, e.g. 5), summaries, and simple names for each side, with each of these generated from the seed dataset (and having no knowledge of the sentences generated thereafter).
4 Evaluation Approach
We compare the novel paired completion method with a total of four comparison approaches, representing a mix of traditional NLP and transformer-based LLM methods – three use a trained logistic regression (Hosmer Jr, Lemeshow, and Sturdivant 2013) classifier over varying training sample sizes, either employing TF-IDF vectors (Sparck Jones 1972; Salton 1983), word embedding vectors (Mikolov et al. 2013), or LLM contextual embeddings (OpenAI 2024b) to represent texts in high dimensional space (see Supplementary Information for details); whilst one uses LLMs via a prompt-based approach (described below).
4.1 LLM Prompting
Starting with a corpus of texts to test, we construct a prompt with three components: 1) a static instructional component which provides the LLM with the task information; 2) a set of context texts that represent framing A and B to be tested ($A,B$); and 3) a single target text ($x$). Unlike in LLM paired completion, we do not require the LLM to provide log-probs for the input sequence, but instead, we obtain the log-probs of the first two tokens produced by the LLM in response to this prompt, i.e. the first two generated tokens. Note that, by virtue of the constraints in the prompt, these probabilities include the log-probs for both response A and B. We extract the probability of the first token of the label assigned to A (e.g. ‘[equality]’ [1 token]), and B (e.g. ‘[mis][og][yny]’ [3 tokens]), respectively. With this information we can both identify which set the LLM has assigned the text to (based on the higher probability of its tokens) and calculate the equivalent Diff metric, $\Delta(A,B,x)$.
We use a fixed prompt, which was initially fine-tuned for GPT-4 and GPT-3.5, and then further tuned for Mixtral-8x7b-Instruct-v0.1 and LLaMA-2-70B-Chat. In hindsight, it was a mistake to tune our prompts for GPT-4 first, as while GPT-4 was very likely going to give the best performance on the tasks at hand (compared to the other models in consideration), it was also a lot more forgiving of errors, confusing wording, and conflicting instructions within the prompt. It is also possible this somewhat biased the prompts towards the OpenAI models, but this was unavoidable given our preliminary results (not included in this paper) were gathered using only the OpenAI APIs; in any case, our results demonstrate the superiority of the open-source models on these tasks. We used a single prompt across all models in our final experiments.
4.2 Performance Analysis
In terms of true-positives (TP), false positives (PF) and false negatives (FN), the F1 score is calculated as,
$f1=\frac{TP}{TP+\frac{1}{2}(FP+FN)}\,\,.$
The f1 score takes a value from 0 to 1, and will be equal to 1 when the method perfectly identifies all the ‘As’ in the data, and does not mis-identify any ‘Bs’ as ‘As’.
Confidence intervals (95%) for f1 scores were either calculated directly from replicates, in the case of the logistic regression methods (TF-IDF, word- and LLM- embeddings), and using bootstrapping in the LLM API cases (100 replicates, 1000 samples).
4.3 Summary of experiments
Together, across the five methods, four topics, and related variants, 192 experiments were conducted, as summarised in Table 1.
Method Models Topics Variants Total
LR:TF-IDF 1 4 6 24
LR:FastText 1 4 6 24
LR:LLM Embed. 2 4 6 48
LLM Paired Compl. 4 4 2 32
LLM Prompt. 4 4 4 64
TOTAL 192
Table 1: Summary of Experiments The same four topics were tested across all configurations (‘dog-ownership’, ‘climate-change’, ‘domestic violence’, ‘misogyny’). For each LR (Logistic Regression) style experiment, 6 different sub-set sizes were used ($n\in\{10,20,50,100,200,500\}$). For LLM Paired Completion two variants for the number of conditioners were used ($k\in\{1,2\}$). For LLM Prompting, 4 prompt variants were used (seeds, distilled, summary, zero-shot).
5 Results & Discussion
Our experiments demonstrate strong performance across the board for both prompt-based and paired completion methods, as shown in Figure 2. Paired completion methods tend to statistically perform the same or better than prompt-based methods. This section includes a broad summary of results. More detailed results, tables, and discussion can be found in the appendices.
Figure 2: F1 outcomes across LLM prompting ($\square$) and paired-completion ($\triangle$). Filled markers indicate approaches that are statistically similar to most performant method. Semi-transparent shading shows 95% confidence interval for these methods to indicate other methods which provide similar performance to performant models. Ranking is by overall performance. See appendix for performance comparison with log-reg classification methods.
5.1 Comparative Analysis of Classification Methods
With sufficient data (200+ samples), the embedding approach was competitive with GPT-4 prompting. However, embeddings performed significantly worse in few-shot learning contexts. Among LLM instruct models, GPT-4-Turbo outperformed all other models. GPT-3.5-Turbo, Mixtral-8x7b-Instruct-v0.1, and LLaMA-2-70B-Chat had similar performance, with LLaMA-2-70B-Chat having the highest propensity for failure modes. For the paired completion approach, performance trended with model parameter count, with LLaMA-2-70B performing best, followed by Mixtral-8x7b, davinci-002, and babbage-002. This consistency may occur because paired completion is less sensitive to model-specific factors like architecture, alignment, and fine-tuning.
The three methods for interaction with LLMs that were analysed are all effective, and uniquely suited to different scenarios. The paired completion approach proved highly effective, efficient, and robust to model-specific influences. Embedding-based methods are extremely cheap due to a combination of cheap models and fewer calls to the APIs, and proved very effective with sufficient data. However, this data threshold was far beyond the five exemplars used for the other two approaches, and performance suffered greatly when using 50 exemplars (which is still an order of magnitude more than the five exemplars provided to the other two methods). Prompt-based completion proved effective, particularly with GPT-4 (which does not support the logit outputs required for paired completion), but when possible we generally found paired completion to be more cost-effective than the prompting approach. On the other hand, the prompt-based method is very straightforward, and as models increase in capability (and, ideally, increase in cost-performance as well) the cost-performance distinction between these two techniques might diminish.
5.2 Cost vs. Performance
Figure 3: Cost — performance trade-off for LLM methods. Colouring and styling follows Fig 2. Model short name and variant are provided for clarity.
An analysis of the cost-performance trade-off for the LLM methods (Figure 3) reveals that the paired completion approach with LLaMA-2-70b and Mixtral-8x7b is very cost-effective for their level of performance. While GPT-4 had the best overall performance, it was also by far the most expensive. Other configurations can be chosen based on requirements and funding availability. All LLM-based approaches were significantly more expensive than the embedding approaches, which require more data but proved competitive given sufficient training examples.
5.3 Model Bias
We observed differences in the bias displayed by models and techniques that were dataset-dependent (Figure 4). Embedding-based approaches appear most robust to bias, with no statistically significant bias found for any configuration. LLM-based approaches demonstrated bias in some scenarios, with the $k=2$ paired completion configuration potentially reducing bias compared to $k=1$. The top performing LLM paired completion methods (mxtrl-k=2; llama-k=2) show significantly less bias than the top LLM prompting approaches, including GPT-4. Further studies are needed to examine the sources of these biases, such as bias in training data, language modeling, or alignment. However, the results suggest the stronger LLM paired completion methods (e.g. llama-k=2) achieve a balance of high accuracy and low bias.
Figure 4: Inference Bias Mean and (95% confidence intervals) for issue-framing asymmetry, or model bias, calculated as the difference between the off-diagonals in a normalised confusion matrix. Scores are given where non-zero bias is statistically significant ($p<0.05$).
6 Limitations & Further Work
Use of synthetic data for evaluation
- Whilst using synthetic data for evaluation has some benefits (described earlier) there are also some significant limitations. Principally, we cannot be sure that our evaluation results would carry over to a human-labelled ground-truth data set. In the ideal case, a large expert annotation activity would be undertaken to generate a conceptual framing dataset covering a range of issues, and dimensions, and conceptual framings. Such a dataset would be of huge benefit to the field and would no doubt spur further refinement of framing analysis methods. As an intermediate step, a representative sample of our synthetic dataset could be validated by expert annotation to provide some comfort to our main findings.
Opposing framings, extension beyond ‘support,oppose’
– Our study leverages a traditional two-class classification paradigm by seeking each method’s determination as to whether a given target text is drawn from synthetic framing set ‘a’ or ‘b’. Whilst parsimonious and a reasonable starting position, a natural question arises as to how paired completion could be extended to any number of framings on a given dimension of an issue. For example, climate change (issue) causation (dimension) could be explored with three framings around: ‘anthropogenic activity’, ‘volcanic activity’, or ‘solar cycles’. The time complexity of the comparison algorithm after calling the model is possibly superlinear in the number of classes (though the exact behaviour would be implementation-dependent), but the number of model calls is linear in the number of classes. We thus expect the latter to be the overwhelming factor at least until reaching thousands of classes, given an efficient implementation of the comparison algorithm.
Model bias in aligned models
– The data seems to offer some support for the conjecture that aligned models are more prone to bias when performing framing alignment, but we cannot make any definitive claims without significantly more evidence and data. We only used three “serious” topics (climate change, domestic violence, and misogyny); for further study, we would significantly expand this (perhaps to 10, 20, or even 100 topics, ranging across and beyond, say, the Overton window (Russell 2006)).